# Deep-Learning-Driven Prefetching for Far Memory

### v1

@@
Abstract.
Modern software systems face increasing runtime performance demands, particularly in emerging architectures like far memory, where local-memory misses incur significant latency. While machine learning (ML) has proven effective in offline systems optimization, its application to high-frequency, runtime-level problems remains limited due to strict performance, generalization, and integration constraints. We present FarSight, a Linux-based far-memory system that leverages deep learning (DL) to efficiently perform accurate data prefetching. FarSight separates application semantics from runtime memory layout, allowing offline-trained DL models to predict access patterns using a compact vocabulary of ordinal possibilities, resolved at runtime through lightweight mapping structures. By combining asynchronous inference, lookahead prediction, and a cache-resident DL model, FarSight achieves high prediction accuracy with low runtime overhead. Our evaluation of FarSight on four data-intensive workloads shows that it outperforms the state-of-the-art far-memory system by up to 3.6 times. Overall, this work demonstrates the feasibility and advantages of applying modern ML techniques to complex, performance-critical software runtime problems.
1. Introduction
Machine learning (ML) has recently demonstrated strong potential in addressing a variety of systems challenges. Prominent examples include optimizing operating system and virtual machine configurations (yang2022stun, ; cheng2024dynamic, ; cano2018adares, ), predicting request arrival times for cluster-level resource management (resource-central, ; ml-datacenter, ; peng2019dl2, ; Paragon, ), and estimating object hotness to improve data placement in heterogeneous memory (ML-Object-HeteroMem, ; lu2024hammerefficienthotcolddata, ; IDT-RL-TieredMem, ). These successes typically rely on models that are either trained and evaluated offline or invoked infrequently at runtime—such as during object allocation. However, the performance of applications is directly influenced by frequent, fine-grained events like memory accesses and instruction execution.
So far, ML-based solutions for such high-frequency events are largely confined to the micro-architectural level. Examples include CPU branch prediction (TinyBERT-BP, ; tarsa2019-bp, ), instruction prefetching (pierce1994wrongpath, ; ramesh2024pdip, ; protean2023, ), and cache line prefetching (duong2024twilight, ; zhang2023dart, ; voyager, )). Unlike hardware-level ML techniques that operate on small, consistent inputs with tight latency and resource constraints, applying ML at the software systems runtime layer introduces distinct challenges and opportunities. At the systems level, application behavior can vary widely by input data, application configurations, and environments, and application performance is impacted by longer-term history. Meanwhile, system software has more flexible access to idle CPU cores and memory space, potentially allowing for more resource-demanding ML techniques. This discrepancy raises a fundamental research question:
How to properly leverage ML techniques to solve complex, performance-critical runtime problems at the systems layer?
To answer this question, we explore a representative and challenging problem: ML-based far memory prefetching in the Linux kernel. Far-memory systems—where most of the data resides in remote memory and only a small subset is cached locally—are gaining traction for reducing overall memory costs through the use of cheaper, network-attached memory and memory sharing across nodes. In far-memory systems, application performance is highly sensitive to the latency of fetching data on demand from remote memory, as a far-memory access over RDMA takes more than 20 times of a local DRAM access.
To make far memory viable, the on-demand data fetching overhead must be mitigated. Today’s predominant approach is to improve far-memory communication bandwidth and latency with interconnect technologies like CXL (cxl_spec2023, ; chen2023pond, ; wang2023dmemos, ; subramaniam2022cxl, ; chen2024nextgen, ). However, these approaches require substantial hardware upgrades and ecosystem adoption, slowing their deployment. An alternative and orthogonal strategy is to reduce the frequency of on-demand fetches through software prefetching. Today’s far-memory systems employ rule-based prefetchers (e.g., sequential or strided access patterns (mira, ; almaruf2020leap, )), which perform well for regular memory access patterns but fail in the face of complex access behaviors influenced by non-trivial interactions between application phases, control flow, data locality, and cache hierarchy.
Many data-center workloads including graph processing (PageRank, ; han2024graph, ), tree and index structures (guttman1984r, ; gusfield1997algorithms, ), pointer chasing (hsieh2017implementing, ), and recursive data structures (harold2004xml, ) exhibit memory access patterns that defy rule-based prefetching. If these access patterns could be learned and predicted accurately, far-memory systems could proactively fetch data and mitigate the performance penalties associated with remote access, even in the absence of new hardware.
Figure 1. FarSight Achieving Three Key Goals Together. FarSight, FastSwap (fastswap, ), and Hermit (hermit, ) are far-memory systems that run in the Linux kernel. Voyager (voyager, ), Hashemi etal. (pmlr-v80-hashemi18a, ), and Twilight (duong2024twilight, ) are micro-architecture CPU cache prefetchers implemented in simulation or with offline traces.
Modern deep-learning (DL) techniques such as LSTM (hochreiter1997long, ), RNN (socher2011parsing, ), and Transformer (vaswani2017attention, ) have demonstrated the ability to model complex, long-range, and nonlinear patterns that traditional heuristics cannot capture. However, incorporating such models in performance-critical system settings introduces a unique set of challenges. To ensure good application performance, ML techniques must impose minimal overhead. This makes online training impractical, as it introduces significant memory and computational costs. Offline training, while avoiding these runtime expenses, raises a different challenge: ensuring the trained model remains effective on new or unseen inputs. At runtime, predictions must be both accurate and lightweight–—incorrect predictions lead to prefetching irrelevant data, which wastes local memory and degrades application performance. Although larger models and longer input histories can enhance prediction accuracy, they also increase the cost of access tracking, inference latency, and memory usage due to the size of model weights. Ultimately, reconciling the competing goals of high accuracy, low overhead, and strong generalization remains a fundamental and difficult challenge (Figure 1).
@@ 1. Introduction
To confront this challenge and demonstrate the benefits of ML techniques for performance-critical systems problems, we build FarSight, a Linux-based far-memory system that incorporates modern DL techniques for effective data prefetching. Our key insight is that memory access behavior is governed by both application semantics (e.g., algorithmic logic) and input-dependent runtime context (e.g., memory layout). The semantics tend to generalize across inputs and can be learned offline, but the actual memory addresses are input-specific and best handled at runtime. FarSight exploits this separation by training a DL model to learn semantic patterns and delegating address resolution to a runtime system component.
@@
Specifically, we propose to represent application semantics as relationships between memory accesses. For each access, we observe that the subsequent access usually only has a small set of possibilities. We assign each possible outcome an ordinal identity. While the actual memory addresses corresponding to these possibilities vary across different inputs, the transition pattern—i.e., which ordinal is likely to follow given the history–—is often learnable and generalizable. For instance, in a linked list traversal, each accessed node is always followed by the next node in the list. Although the concrete addresses of these nodes differ per execution, the access behavior remains the same. Similarly, in the PageRank algorithm, a page’s rank update depends on the ranks of linking pages. The specific memory addresses of those linked pages depend on the input graph, but their identities can be precisely captured at runtime and predicted using ordinal labels.
@@
To formalize the above idea, we set the DL model vocabulary as the anticpated outcome possibilities (i.e., a configurable $K$ defaulting to 64). The DL model uses memory access history sequences encoded into $K$ vocabulary and predicts future memory accesses as a sequence of 0 to $K-1$ ordinals. These ordinals are resolved at runtime via a lightweight future map–—a data structure that records actual addresses observed during execution. This design significantly reduces model vocabulary from the full memory address space to a small, fixed size, enabling high prediction accuracy with a compact DL model.
Building on top of this idea, we optimize FarSight’s performance while delivering high prediction accuracy with several techniques. (1) Asynchronous prediction and prefetching: Prediction is triggered when application threads block on far-memory fetches, utilizing otherwise idle CPU cycles and hiding inference latency behind the critical path. Afterward, prefetching requests are executed asynchronously in the background. (2) Multi-step lookahead: The model predicts several steps ahead, enabling timely prefetching before the application issues the corresponding memory accesses. (3) Compact DL architecture: We employ Retentive Network (retnet, ), a compact RNN-based model small enough to fit within a core’s L1 cache, ensuring low latency and energy efficiency. (4) Efficient input encoding: We use a position encoding scheme that supports reuse of cached context across predictions, reducing redundant computation and memory access overhead.
@@
We implement FarSight’s training as an offline process when an application is deployed. We implement FarSight’s runtime system in the Linux kernel, with most of it being a Linux kernel module and the rest changing Linux’s swap system slightly. We evaluate FarSight with four real-world applications and benchmarks: MCF (spec2006_429mcf, ), PageRank and Shortest Path from the GAP benchmark suite (beamer2015gap, ), and XGBoost (chen2016xgboost, ). We compare FarSight to FastSwap (fastswap, ) and Hermit (hermit, ), two Linux swap-based far memory systems. Our results show that FarSight outperforms FastSwap by up to 3.6 times and Hermit by up to 2.6 times.
Overall, this paper makes the following key contributions.
  • The first ML-based far-memory system, FarSight.
  • Three insights of opportunities and challenges in ML-based prefetching.
  • The novel idea of decoupling memory-access semantics and memory address layouts.
  • Full implementation of a DL-based swap system in the Linux kernel.
  • Various optimizations at the systems and prediction-method layers.
We will open source FarSight upon the paper’s acceptance.
2. Motivation and Related Works
This section first provides an overview of far-memory systems and far-memory prefetching. We then discuss why and when ML techniques are a good fit for far-memory prefetching. Finally, we discuss the challenges of applying DL techniques for far-memory prefetching.
2.1. Far Memory and Existing Prefetch Approaches
@@
Far-memory systems are systems where applications have access to memory beyond CPU-local memory, e.g., memory at another server or memory in a disaggregated memory pool. Far memory allows applications to access larger amounts of memory and more efficient cluster memory utilization. For applications to utilize far memory, there usually is an indirection layer (e.g., a swap system (fastswap, ; hermit, ) or a user-space library (aifm-osdi20, ; mira, )) that fetches data from far to local memory.
The main limitation of far-memory systems is the communication delay between local and far memory. For example, for a local memory size that is half of the far memory size, naive implementation of a far-memory system could result in half of the accesses going to far memory, resulting in an application slowdown of 13 times for RDMA-based far-memory systems. In practice, to strive for higher memory resource efficiency, local memory sizes are often set to below half of the far memory size (chen2023pond, ).
@@
To hide this delay, most far-memory systems prefetch future accesses from far memory and cache them locally. Existing far-memory systems (fastswap, ; aifm-osdi20, ; almaruf2020leap, ) prefetch far-memory data with rule-based approaches by detecting and following linear and strided patterns. As such, they are limited to only benefit applications with such regular memory access patterns.
2.2. Why and When ML for Far-Memory Prefetching
A significant number of data-intensive applications have memory access patterns that simple rules cannot capture. For example, graph-based algorithms like PageRank (PageRank, ) and graph search have access patterns that are highly dependent on graph structures. As another example, ML algorithms have memory access patterns dependent on ML model architectures. Figure 2 visualizes memory accesses in PageRank using the Twitter graph with 41M nodes (snap-twitter2010, ) and insertion sort of 1 million randomly generated values in a linked list. As seen visually, these applications follow specific memory-access patterns, but they are hard to describe by simple rules. This is because these applications have repeatable program logic that is not completely random or completely rule-based.
More generally, we expect ML techniques to work for applications with code pieces that are repeatedly executed (e.g., a loop, a recursive function, etc.) or follow some traversal behavior (e.g., link-list walk, graph walk, indirect array accesses, array accesses according to some algorithm, etc.). Many data-center applications, such as data analytics, ML algorithms, graph processing, tree-based indexes, sorting, etc., fit these features. On the other hand, memory accesses that are completely random or follow simple rules do not fit ML-based approaches; for these applications, FarSight falls back to using rule-based prefetching or no prefetching and thus performs similarly to existing far-memory systems.
Figure 2. Example Memory Access Sequence over Time Captured by tracing page access sequence during execution.
Insight 1. Many data-center applications show repeatable, patterned memory-access behaviors that can potentially be captured by ML but not by simple rules.
2.3. Challenges of DL for Far Memory Prefetching
@@
Performance and why not GPU. Unlike many existing exploration of ML for systems, prefetches are on the performance-critical paths that directly impact application performance. Prefetched data that arrive later than the on-demand accesses make them useless. Pausing applications to perform prediction and prefetching is also not an option, especially when they take long.
To put things in perspective, we measure model prediction time (also known as model forwarding time) on GPUs. Predicting one token with 64 input tokens using the Llama-7B model on an Nvidia A100 GPU takes around 15 milliseconds; even small models like NanoGPT-10M take hundreds of microseconds. To understand where the latency comes from, we build a minimal “model” with only one matrix multiplication and memory copying from/to the CPU. With this, we find that the launching time for the matrix multiplication and memory copy kernels takes close to 10 microseconds. In comparison, a far-memory 4 KB page read takes around 2 microseconds with today’s InfiniBand-based network.
The long delay of GPU-side model forwarding implies that the model needs to predict far into the future. This is because the memory access history that the model uses is before the forwarding starts, and by the time forwarding finishes, hundreds of thousands of memory accesses could have happened. Effective prefetching requires a prediction of at least this many accesses in the future, but long predictions are usually less accurate.
Insight 2. GPU-based memory-access prediction is slow and thus requires predicting farther into the future, making accurate prediction hard.
Accuracy. Although wrong prefetches do not affect application execution correctness, they waste local memory space and network bandwidth, which are especially precious under far-memory environments. As local memory is expected to run to capacity, a prefetched page will need another local page to be swapped out on demand, taking about 4 microseconds from our evaluation. Thus, it is essential to design ML techniques for accuracy. Model accuracy depends not just on model size but also largely on the prediction problem formation. For example, longer input history usually helps model prediction to be more accurate, but at the expense of higher monitoring overhead. As another example, a small vocabulary (i.e., number of distinct prediction values) is easier for a model to predict accurately, but makes it challenging to represent a complex problem space like memory addressing.
@@
Generalization. Because of the performance requirement, it is infeasible to train or fine tune a model during the run time. An offline trained model avoids any runtime performance overhead but has no access to runtime status. Although many data-center applications follow memory-access patterns, the exact pattern is often input-dependent. Moreover, the actual memory addresses being accessed can be different across runs even with the same input because of memory address randomization techniques like ASLR.
Insight 3. It is challenging to achieve performance, accuracy, and generalization together, as they have competing goals.
3. FarSight Design
FarSight is a swap-based far-memory system that prefetches memory pages from network-attached far memory with an DL-based memory access predictor. FarSight consists of an offline training component and an online predictor and prefetcher component sitting in the Linux kernel’s swap system, as illustrated in Figure 3. We choose to build FarSight as a Linux-kernel-based swap system, as it allows all Linux-compatible applications to transparently use far memory (fastswap, ; hermit, ; canvas, ; almaruf2020leap, ). The core of FarSight can also be implemented in the user space.
@@ 3. FarSight Design
When an application is deployed, FarSight trains a small model (3K-parameter Retentive Network (retnet, ) architecture) by tracking the execution of the application with user-supplied sample application inputs. During the run time, FarSight loads the trained model onto each CPU core running the application. FarSight predicts future far-memory accesses using its captured recent program execution history and issues the corresponding prefetch requests. When the prefetch is correct, future access will be local. Otherwise, or when prefetch is slower than the future access, the application thread issues an on-demand far-memory access.
This section discusses the core ideas and technical details of FarSight.
3.1. Two Key Ideas
Based on our insights from §2, we propose two primary ideas for FarSight.
Figure 3. FarSight Overall Architecture All red parts are FarSight.
Pattern and addressing decoupling. The first core idea revolves around how we frame the prediction task. We observe that an application’s memory access patterns often exhibit repetition due to constructs like loops and recursion. However, these patterns are typically complex and nonlinear, making them difficult to model with simple rules or heuristics. DL models, on the other hand, are well-suited to capturing such long-range, nonlinear dependencies. Nevertheless, to maintain efficiency—particularly with smaller models that can run within a single CPU core and are trained offline—we must simplify the problem space and minimize runtime variability.
Our idea is to decouple application memory access semantics from the actual runtime memory layout by using DL prediction for the former and mapping tables for the latter. Specifically, we use a small DL model to predict memory access relationships in terms of abstract ordinals—–representing the possible memory-access outcomes after a short history of memory behavior—–rather than concrete memory addresses or offsets, which are highly input- and environment-dependent. At runtime, we construct future maps: mapping tables that resolve these predicted ordinals to actual memory addresses observed during the program’s first access, thus capturing the true memory layout dynamically.
In contrast to prior approaches that rely exclusively on ML prediction (§5) or purely on runtime history (as in traditional rule-based prefetching), FarSight integrates both DL prediction and runtime recording. By structurally decoupling the problem space, each component operates where it is most effective.
CPU prediction and I/O overlapping. Our second main idea focuses on the efficient realization of the DL-based prediction framework. As discussed in §2, GPU-based prediction has performance and accuracy issues. Thus, FarSight performs its model prediction in CPU. To avoid the energy and performance cost of additional CPU cores, FarSight performs prediction for an application on the same core it runs on. Our idea to avoid application performance overhead is to hide model prediction behind far-memory I/O time. Specifically, FarSight performs prediction when a foreground page fault is being handled with far-memory data read. We then issue asynchronous prefetch requests in the background.
@@ 3.1. Two Key Ideas
Figure 4. FarSight Prediction Representation An example of vocabulary size ($K$) being 4. The top part shows code/algorithm corresponding to the accesses of chunks addr-x, addr-b, and addr-e. The bottom shows the input to the model: the chunk addresses and PCs of the 5 previous misses.
3.2. Prediction Task Formation

### v2

# Learning Semantics, Not Addresses: Runtime Neural Prefetching for Far Memory

@@
Abstract
Memory prefetching has long boosted CPU caches and is increasingly vital for far-memory systems, where large portions of memory are offloaded to cheaper, remote tiers. While effective prefetching requires accurate prediction of future accesses, prior ML approaches have been limited to simulation or small-scale hardware. We introduce FarSight, the first Linux-based far-memory system to leverage deep learning by decoupling application semantics from runtime memory layout. This separation enables offline-trained models to predict access patterns over a compact ordinal vocabulary, which are resolved at runtime through lightweight mappings. Across four data-intensive workloads, FarSight delivers up to 3.6× higher performance than the state-of-the-art.
1 Introduction
In response to increasing application memory demands and the slowing down of server main memory scaling, major datacenters like Google and Microsoft have adopted the datacenter architecture of far memory, where most of the data resides in remote, network-attached memory and only a small subset is cached locally in CPU-attached DRAM or GPU HBM Chen et al. (2023); Lu et al. (2024); Liu et al. (2025). Far-memory systems offer lower cost/GB, higher energy efficiency, lower carbon footprint, and elastic capacity. However, the main limitation of today’s far-memory systems is the application performance overhead caused by accessing non-local data from remote memory, an operation that is often more than 20 times slower than a local DRAM access. A common mechanism employed in existing far-memory systems is to fetch data from far memory in the background before they are accessed (i.e., prefetching), thereby avoiding the foreground, on-demand far-memory data fetching time that directly affects application performance.
Effective prefetching relies on the accurate prediction of future memory accesses. Today’s far-memory systems take a conservative approach of only prefetching accesses that follow simple rules (e.g., sequential or strided access patterns Guo et al. (2023); Maruf & Chowdhury (2020)). However, most data-center workloads, such as graph processing Lawrence et al. (1998); Han et al. (2024), tree and index structures Guttman (1984); Gusfield (1997), pointer chasing Hsieh et al. (2017), and recursive data structures Harold & Means (2004), exhibit memory access patterns that defy rule-based prefetching. As such, they still suffer from huge performance overheads, preventing the wide adoption of far-memory systems and their potential cost saving which datacenters could have otherwise achieved.
We propose FarSight, a deep-learning-based prefetching mechanism designed for far-memory systems and implemented in the Linux kernel. Different from simulation- or hardware-based CPU cache prefetchers, a key challenge in software-based, far-memory prefetchers is to avoid prediction latency overhead while achieving high accuracy in a vast search space. For example, any online training or even traveling through the PCIe bus to a GPU for model inference would significantly slow down application execution. Representing memory accesses directly using their addresses in a model is also infeasible (a 64-bit machine has $2^{48}$ virtual memory addresses).
@@ 1 Introduction
Our insight is that memory access behavior is governed by both application semantics (e.g., algorithmic logic) and input-dependent runtime context (e.g., memory layout). The semantics tend to generalize across inputs and can be learned offline, but the actual memory addresses are input-specific and best handled at runtime. We exploit this separation by training a deep-learning (DL) model to learn semantic patterns and delegating address resolution to a runtime system component.
@@
Specifically, we propose to represent application semantics as relationships between memory accesses. For each access, we observe that the subsequent access usually only has a small set of possibilities. We assign each possible outcome an ordinal. While the actual memory addresses corresponding to these possibilities vary across different inputs, the transition pattern—i.e., which ordinal is likely to follow given the history–—is often learnable and generalizable. For instance, in a linked list traversal, each accessed node is always followed by the next node in the list. Although the memory addresses of these nodes differ per execution, the access behavior remains the same.
@@
We set the DL model vocabulary as the anticipated outcome possibilities (i.e., a configurable $K$ defaulting to 64). The DL model uses memory access history sequences encoded in the vocabulary $K$ and predicts future memory accesses as a sequence of ordinals ranging from 0 to K1K-1. These ordinals are resolved at runtime via a lightweight future map–—an in-memory mapping table we propose to record for each accessed memory. A future map contains $K$ entries, each mapping (from its index) to a memory address observed at runtime to follow the access to the current page. This design significantly reduces model vocabulary from the full memory address space to a small, fixed size, enabling high prediction accuracy with a compact DL model.
Our model builds on top of Retentive Network Sun et al. (2023), a compact Transformer-variant model small enough to fit within a CPU core’s L1 cache. Apart from the above representation, we propose several techniques to further reduce or hide the model prediction and prefetching latencies. We use a position encoding scheme that supports reuse of cached context across predictions, reducing redundant computation and memory access overhead. We predict several accesses ahead so that prefetched data can arrive at the local memory before being accessed, and we perform the prediction and prefetching in the background to hide their latencies.
@@
We implement FarSight’s training as an offline process when an application is deployed. We implement FarSight’s runtime system in the Linux kernel, with most of it being a Linux kernel module and the rest changing Linux’s swap system slightly. We evaluate FarSight with four real-world applications and benchmarks: MCF on SPEC 2006 Benchmark  Henning (2006), PageRank and Shortest Path from the GAP benchmark suite Beamer et al. (2015), and XGBoost Chen & Guestrin (2016). We compare FarSight to a SoTA DL-based memory prefetcher, Twilight Duong et al. (2024), and two non-ML-based far memory systems, FastSwap Amaro et al. (2020), Hermit Qiao et al. (2023). Our results show that FarSight outperforms FastSwap by up to 3.6 times, Hermit by up to 2.6 times, and Twilight by up to 3.9 times.
2 Background, Related Works, and Motivation
2.1 Far Memory and Existing Prefetch Approaches
@@
Far-memory systems are systems where applications have access to memory beyond CPU-local memory, e.g., memory at another server or memory in a disaggregated memory pool. Far memory allows applications to access larger amounts of cheaper (e.g., older generations of, non-volatile, pooled) memory. In practice, to strive for higher memory resource efficiency, local memory sizes are often set to below half of the far memory size Chen et al. (2023). For applications to utilize far memory, there usually is an indirection layer (e.g., a swap system Amaro et al. (2020); Qiao et al. (2023) or a user-space library Ruan et al. (2020); Guo et al. (2023)) that fetches data from far to local memory.
@@
The main limitation of far-memory systems is the communication delay between local and far memory. For example, for a local memory size that is half of the far memory size, naive implementation of a far-memory system could result in half of the accesses going to far memory, resulting in an application slowdown of 13 times for RDMA-based far-memory systems. To hide this delay, most far-memory systems prefetch future accesses from far memory and cache them locally. Existing far-memory systems Amaro et al. (2020); Ruan et al. (2020); Maruf & Chowdhury (2020) prefetch far-memory data with rule-based approaches by detecting and following linear and strided patterns. As such, they are limited to only benefit applications with such regular memory access patterns.
Prior research works Shi et al. (2021); Hashemi et al. (2018); Srivastava et al. (2019); Peled et al. (2015; 2019) explored using ML techniques for local server prefetch predictions at the micro-architecture level by prefetching data from memory into CPU cache. However, because of their performance and/or accuracy issues, they have only been realized in simulation or for offline trace analysis. For example,  Peled et al. (2015) proposed reinforcement-learning-based and regression-based Peled et al. (2019) approaches for memory prefetching. The former sets up the prefetching prediction as a classification problem and can only accommodate four possible address offsets for each prediction. The latter has accuracy issues as a regression model aims to be close to the ground truth, but correct prefetch requires the exact truth. Twilight and T-LITE Duong et al. (2024) use the combination of a customized two-layer neural-network model, clustering, and frequency-based history table for CPU cache prefetching. DART Zhang et al. (2023) distills a transformer model and then transforms the distilled model into a hierarchy of table lookups to reduce runtime performance overhead. Although these works have shown their CPU cache prefetch effectiveness through simulation, their training and prediction processes are complex and lack generalization or consistent accuracy.
2.2 Challenges of DL for Far Memory Prefetching
@@
[…]
@@
Generalization. Because of the performance requirement, it is infeasible to train or fine tune a model at run time. An offline-trained model avoids any runtime overhead but has no access to runtime status. An application’s memory accesses at runtime are often input-dependent. Moreover, memory addresses can be different across runs even with the same input because of memory address randomization techniques like address space layout randomization (ASLR).
To make DL-based prefetch practical for far-memory systems, latency, accuracy, and generalization must be achieved simultaneously.
Figure 1: FarSight overall architecture all red parts are FarSight.
3 FarSight Design
@@
FarSight consists of an offline training component and an online predictor and prefetcher component sitting in the Linux kernel’s swap system, as illustrated in Figure 1. When an application is deployed, FarSight trains a small model (3K-parameter Retentive Network Sun et al. (2023) architecture) by tracking the execution of the application with user-supplied sample application inputs. During runtime, FarSight loads the trained model onto each CPU core running the application. FarSight predicts future far-memory accesses using its captured recent program execution history and issues the corresponding prefetch requests.
@@
Figure 2: FarSight prediction representation An example of vocabulary size ($K$) being 4. The top part shows code/algorithm corresponding to the accesses of chunks addr-x, addr-b, and addr-e. The bottom shows the input to the model: the chunk addresses and PCs of the 5 previous misses.
Figure 3: FarSight’s prediction optimization methods Demonstrating the use of each history window to predict $s$ misses ahead of time and predicting $f=2$ pages at a time.
3.1 Prediction Task Formation